how do i get the points to graph y=2x(squared)+11x+5?
Answer by stanbon(75887)

get the points to graph y=2x(squared)+11x+5?
----
If you know the graph is a parabola, you
only need 3 points.
y = 2x^2 + 11x + 5
---
y-intercept: (0,5)
---
Find the x-intercepts:
2x^2 + 11x + 5 = 0
2x^2 +10x + x + 5 = 0
Factor:
2x(x+5) + (x+5) = 0
(x+5)(2x+1) = 0
x = -5, or x = -1/2
--------
Points: (-5,0) and (-1/2,0)
---
Plot those three points and draw a parabola opening upward thru them.
=====================
Cheers,
Stan H.
=====================